OpenStudy (anonymous):

hey is there anyone who can help me with converting a triple integral to cylindrical coordinates?? cheers!

OpenStudy (anonymous):

particularly with the example...: I=int[-1 to 1] int[-sqrt(1-x^2) to sqrt(1-x^2)] int[0 to 2*sqrt(x^2 + y^2)] x^2dzdydx

OpenStudy (anonymous):

answer is given as \[\int\limits_{0}^{2\Pi}\int\limits_{0}^{1}\int\limits_{0}^{2r}r^2\cos^2\Theta rdzdrd \Theta\] and I understand everything apart from how to get 0 to 2Pi for the theta coordinates.... Help please! :D

OpenStudy (anonymous):

Sorry - original equation is \[\int\limits_{-1}^{1}\int\limits_{-\sqrt(1-x^2)}^{\sqrt(1-x^2}\int\limits_{0}^{2\sqrt(x^2+y^2)}x^2dzdydx\]

OpenStudy (anonymous):

Feldy, it's a circle. The limits of your dy part of your integral are y = sqrt(1-x^2) and y = - sqrt(1-x^2). If you square both sides on each of these equations, you will get the equation of a circle with radius one. If you draw the cricle, draw a line vertically simple since dy is first, it touches the top and the bottom of the circle. Therefore, the region of intrest spans the entire circle....from 0 to 2 pi.

OpenStudy (anonymous):

your volume is the area between a cone and the xy plane over the unit disc.

OpenStudy (anonymous):

x^2+y^2 = r^2 for the polar substution

OpenStudy (anonymous):

x= r cos(theta)

OpenStudy (anonymous):

and the jacobian of the substitution is r

OpenStudy (anonymous):

the limits on x and y give a unit disc which in polar coordinates is 0<=theta<=2pi and 0<=r<=1

OpenStudy (anonymous):

Cheers guys... so Nisha, if , say my y limits were 0 to y=sqrt(1-x^2), instead of -sqrt(1-x^2), then it would be 0 to Pi, instead of 2Pi???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would only trace out the first and second quadrants

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much!